OpenStudy (shylovesall):

can someone help me understand my one math test I failed and need to do grade recovery?

OpenStudy (kamibug):

Do you have a specific question? :)

OpenStudy (shylovesall):

I have a lot of questions. im not good at math at all.

OpenStudy (kamibug):

Alright, what's your first one?

OpenStudy (shylovesall):

(-4.7)^0

OpenStudy (kamibug):

Any number to the zero power is 1, and zero to any power is 0

OpenStudy (shylovesall):

I don't understand still.. I'm sorry I'm bad at math.

OpenStudy (kamibug):

It's okay. :) Notice that the base number "-4.7" is has an exponent of 0. \[-4.7^0\] This is also called "being raised to the power of zero," like you'd say 2^3 is two raised to the power of 3, or the 3rd power. When you have ANY number at all under the exponent of 0, it is equal to one. For example \[6^0 = 1\] Does that help a bit? :)

OpenStudy (shylovesall):

yes that makes a lot more sense. :) can you help me with more?

OpenStudy (kamibug):

Sure :)

OpenStudy (shylovesall):

(k^-2)^9

OpenStudy (kamibug):

http://www.mathplanet.com/education/algebra-1/exponents-and-exponential-functions/properties-of-exponents ^The above is a good site that'll help you with lots of properties of exponents. :) This this case, we are raising a power to a power. That can be done by *multiplying* each exponent. :D For example \[(z ^{-3})^5 = z ^{-3*5}=z ^{-15}\] So what would be the answer to your problem? :)

OpenStudy (shylovesall):

k^18?

OpenStudy (kamibug):

Close! Remember the 2 is negative. :) So it'd be k^-18.

OpenStudy (shylovesall):

my teacher said that it was 1/k^18

OpenStudy (kamibug):

Oh, I see. That is correct also. Because to take off the negative you would change the term to its reciprocal. :)\[k ^{-18}=\frac{ 1 }{ k ^{18} }\]

OpenStudy (shylovesall):

that makes a lot of sense. my next one is (6q^6)^-4 my teacher said it is 1/1296q^24

OpenStudy (hwyl):

all other numbers but zero with an exponent of zero is equal to one. \(n^0 = 1; \) if \(n \neq 0 \) we follow the order of operation in any case that an exponent is raise to another exponent

OpenStudy (kamibug):

Oh, yes I forgot to mention the case with zero. :D Thanks @hwyl ! \[0^0 = undefined\]

OpenStudy (shylovesall):

okay but how does that equation equal that? I tried and I don't understand

OpenStudy (hwyl):

any negative exponent is a fractional value \(\large n^{-a} = \dfrac{1}{n^{a}} \)

OpenStudy (shylovesall):

oh okay. can you help me with some more?

OpenStudy (mathmale):

I'd be glad to help you. I do ask that you create new post each time, instead of piggy-backing on old ones, and that you share your own work, your own thinking, before asking for help.

OpenStudy (shylovesall):

I only have 3 more questions. can I just ask them all on here instead of posting a new one for each of them?

OpenStudy (hwyl):

BEFORE WE PROCEED, I NOTICED THAT YOU JUST NEED TO GET YOURSELF FAMILIARIZED WITH THE "LAWS OF EXPONENTS" https://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (hwyl):

THE SUPPOSED "LAWS" ARE JUST A MATHEMATICAL OPERATION PROGRESSION AND THERE ARE NO MYSTERIES IN THEM.

OpenStudy (shylovesall):

I think I figured out the rest of them

OpenStudy (hwyl):

just as you would do \(4+4+4 = 4 \times 3 = 12 \) the same progression can be true in the case of \(4 \times 4 \times 4 = 4^{3} = 64\)